# Andreas Petrou-Zeniou

### v3

@@ B.1 Assumptions
By way of contradiction, let us suppose that there exists a sequence $P_{n_{s}}$ such that:
$Pr_{P_{n_{s}}}\left(L^{n_{s}}\leq\Delta^{n_{s}}(P_{n_{s}})\right)\to\eta<1-\beta/2$ (31)
Let us notice that $L^{n_{s}}\leq\Delta^{n_{s}}(P_{n_{s}})$ if and only if, for any $j$ and $j^{\prime}$ in $\{1,\ldots,p\}$, we have:
$\sqrt{\frac{n_{s}}{\hat{\text{var}}^{n_{s}}_{j^{\prime}j}}}\left(\tilde{S}_{X,j^{\prime}}^{n_{s}}-\mu_{X,j^{\prime}}(P_{n_{s}})-\left(\tilde{S}_{X,j}^{n_{s}}-\mu_{X,j}(P_{n_{s}})\right)\right)\leq d_{1-\beta/2}(\hat{\Sigma}^{n_{s}})$
We have that:
$\left|Pr_{P_{n_{s}}}\left(\bigcap_{1\leq j,j^{\prime}\leq p}\sqrt{\frac{n_{s}}{\hat{\text{var}}_{j^{\prime},j}}}\left(\tilde{S}_{X,j^{\prime}}^{n_{s}}-\mu_{X,j^{\prime}}(P_{n_{s}})-\left(\tilde{S}_{X,j}^{n_{s}}-\mu_{X,j}(P_{n_{s}})\right)\right)\leq d_{1-\beta/2}(\hat{\Sigma}^{n_{s}})\right)\right.$
$\quad\left.-Pr_{\xi_{X}\sim\Phi_{\Sigma(P_{n_{s}})}}\left(\bigcap_{1\leq j,j^{\prime}\leq p}\frac{\left(\xi_{X,j^{\prime}}-\xi_{X,j}\right)}{\sqrt{\text{var}_{j^{\prime},j}(P_{n_{s}})}}\leq d_{1-\beta/2}(\Sigma(P_{n_{s}}))\right)\right|$
$\leq\left|Pr_{P_{n_{s}}}\left(\bigcap_{1\leq j,j^{\prime}\leq p}\sqrt{\frac{n_{s}}{\hat{\text{var}}_{j^{\prime},j}}}\left(\tilde{S}_{X,j^{\prime}}^{n_{s}}-\mu_{X,j^{\prime}}(P_{n_{s}})-\left(\tilde{S}_{X,j}^{n_{s}}-\mu_{X,j}(P_{n_{s}})\right)\right)\leq d_{1-\beta/2}(\hat{\Sigma}^{n_{s}})\right)\right.$
$\quad\quad\left.-Pr_{\xi_{X}\sim\Phi_{\Sigma(P_{n_{s}})}}\left(\bigcap_{1\leq j,j^{\prime}\leq p}\frac{\left(\xi_{X,j^{\prime}}-\xi_{X,j}\right)}{\sqrt{\hat{\text{var}}_{j^{\prime},j}}}\leq d_{1-\beta/2}(\hat{\Sigma}^{n_{s}})\right)\right|$
$\quad+\left|Pr_{\xi_{X}\sim\Phi_{\Sigma(P_{n_{s}})}}\left(\bigcap_{1\leq j,j^{\prime}\leq p}\frac{\left(\xi_{X,j^{\prime}}-\xi_{X,j}\right)}{\sqrt{\hat{\text{var}}_{j^{\prime},j}}}\leq d_{1-\beta}(\hat{\Sigma}^{n_{s}})\right)\right.$
$\quad\quad\left.-Pr_{\xi_{X}\sim\Phi_{\Sigma(P_{n_{s}})}}\left(\bigcap_{1\leq j,j^{\prime}\leq p}\frac{\left(\xi_{X,j^{\prime}}-\xi_{X,j}\right)}{\sqrt{\text{var}_{j^{\prime},j}(P_{n_{s}})}}\leq d_{1-\beta/2}(\Sigma(P_{n_{s}}))\right)\right|$
$=o(1)+\left|Pr_{\xi_{X}\sim\Phi_{\Sigma(P_{n_{s}})}}\left(\bigcap_{1\leq j,j^{\prime}\leq p}\frac{\left(\xi_{X,j^{\prime}}-\xi_{X,j}\right)}{\sqrt{\hat{\text{var}}_{j^{\prime},j}}}\leq d_{1-\beta/2}(\hat{\Sigma}^{n_{s}})\right)\right.$
$\quad\quad\left.-Pr_{\xi_{X}\sim\Phi_{\Sigma(P_{n_{s}})}}\left(\bigcap_{1\leq j,j^{\prime}\leq p}\frac{\left(\xi_{X,j^{\prime}}-\xi_{X,j}\right)}{\sqrt{\text{var}_{j^{\prime},j}(P_{n_{s}})}}\leq d_{1-\beta/2}(\Sigma(P_{n_{s}}))\right)\right|$
where the first inequality follows by the triangle and the second equality holds by lemma B.1. We want to show that:
$\left|Pr_{\xi_{X}\sim\Phi_{\Sigma(P_{n_{s}})}}\left(\bigcap_{1\leq j,j^{\prime}\leq p}\frac{\left(\xi_{X,j^{\prime}}-\xi_{X,j}\right)}{\sqrt{\hat{\text{var}}_{j^{\prime},j}}}\leq d_{1-\beta/2}(\hat{\Sigma}^{n_{s}})\right)\right.$
$\quad\left.-Pr_{\xi_{X}\sim\Phi_{\Sigma(P_{n_{s}})}}\left(\bigcap_{1\leq j,j^{\prime}\leq p}\frac{\left(\xi_{X,j^{\prime}}-\xi_{X,j}\right)}{\sqrt{\text{var}_{j^{\prime},j}(P_{n_{s}})}}\leq d_{1-\beta/2}(\Sigma(P_{n_{s}}))\right)\right|=o(1)$ (32)
Let $a_{j^{\prime},j}^{n_{s}}:=\sqrt{\frac{\text{var}_{j^{\prime},j}(P_{n_{s}})}{\hat{\text{var}}_{j^{\prime},j}}}\frac{d_{1-\beta/2}(\Sigma(P_{n_{s}}))}{d_{1-\beta/2}(\hat{\Sigma}^{n_{s}})}$. By assumption B.2 and the continuity of $d$, we have that, for any $\delta>0$, that:
$Pr_{P_{n_{s}}}\left(\left|a_{j^{\prime},j}^{n_{s}}-1\right|>\delta\quad\text{for any}\quad j,j^{\prime}\right)=o(1)$
Similarly, we have that 32 can be rewritten as:
$\left|Pr_{\xi_{X}\sim\Phi_{\Sigma(P_{n_{s}})}}\left(\bigcap_{1\leq j,j^{\prime}\leq p}a_{j^{\prime},j}^{n_{s}}\frac{\left(\xi_{X,j^{\prime}}-\xi_{X,j}\right)}{\sqrt{\text{var}_{j^{\prime},j}(P_{n_{s}})}}\leq d_{1-\beta/2}(\Sigma(P_{n_{s}}))\right)\right.$
$\quad\left.-Pr_{\xi_{X}\sim\Phi_{\Sigma(P_{n_{s}})}}\left(\bigcap_{1\leq j,j^{\prime}\leq p}\frac{\left(\xi_{X,j^{\prime}}-\xi_{X,j}\right)}{\sqrt{\text{var}_{j^{\prime},j}(P_{n_{s}})}}\leq d_{1-\beta/2}(\Sigma(P_{n_{s}}))\right)\right|=o(1)$
To simplify notation, we will take $s:=n_{s}$, and take a further subsequence $s_{k}$. We construct this further subsequence as follows. For each $s$, we choose any sequence of $\delta_{s}>1$ satisfying:
$\left|Pr_{\xi_{X}\sim\Phi_{\Sigma(P_{s})}}\left(\bigcap_{1\leq j,j^{\prime}\leq p}(1\pm\delta_{s})\frac{\left(\xi_{X,j^{\prime}}-\xi_{X,j}\right)}{\sqrt{\text{var}_{j^{\prime},j}(P_{s})}}\leq d_{1-\beta/2}(\Sigma(P_{s}))\right)\right.$
$\quad\left.-Pr_{\xi_{X}\sim\Phi_{\Sigma(P_{s})}}\left(\bigcap_{1\leq j,j^{\prime}\leq p}\frac{\left(\xi_{X,j^{\prime}}-\xi_{X,j}\right)}{\sqrt{\text{var}_{j^{\prime},j}(P_{s})}}\leq d_{1-\beta/2}(\Sigma(P_{s}))\right)\right|=o(1)$
We choose a subsequence $s_{k}$ such that:
$Pr_{P_{s_{k}}}\left(\left|a_{j^{\prime},j}^{s_{k}}-1\right|>\delta_{s_{k}}\quad\text{for any}\quad j,j^{\prime}\right)=o(1)$
It follows that on this further subsequence, we have:
$\left|Pr_{P_{s_{k}}}\left(\bigcap_{1\leq j,j^{\prime}\leq p}\sqrt{\frac{s_{k}}{\hat{\text{var}}_{j^{\prime},j}}}\left(\tilde{S}_{X,j^{\prime}}^{s_{k}}-\mu_{X,j^{\prime}}(P_{s_{k}})-\left(\tilde{S}_{X,j}^{s_{k}}-\mu_{X,j}(P_{s_{k}})\right)\right)\leq d_{1-\beta/2}(\hat{\Sigma}^{s_{k}})\right)\right.$
$\quad\left.-Pr_{\xi_{X}\sim\Phi_{\Sigma(P_{s_{k}})}}\left(\bigcap_{1\leq j,j^{\prime}\leq p}\frac{\left(\xi_{X,j^{\prime}}-\xi_{X,j}\right)}{\sqrt{\text{var}_{j^{\prime},j}(P_{s_{k}})}}\leq d_{1-\beta/2}(\Sigma(P_{s_{k}}))\right)\right|=o(1)$
Of course, we have that:
$\limsup_{k\to\infty}Pr_{\xi_{X}\sim\Phi_{\Sigma(P_{s_{k}})}}\left(\bigcap_{1\leq j,j^{\prime}\leq p}\frac{\left(\xi_{X,j^{\prime}}-\xi_{X,j}\right)}{\sqrt{\text{var}_{j^{\prime},j}(P_{s_{k}})}}\leq d_{1-\beta/2}(\Sigma(P_{s_{k}}))\right)\geq 1-\beta/2$
Implying that on the further subsequence:
$\limsup_{k\to\infty}Pr_{P_{s_{k}}}\left(\bigcap_{1\leq j,j^{\prime}\leq p}\sqrt{\frac{s_{k}}{\hat{\text{var}}_{j^{\prime},j}}}\left(\tilde{S}_{X,j^{\prime}}^{s_{k}}-\mu_{X,j^{\prime}}(P_{s_{k}})-\left(\tilde{S}_{X,j}^{s_{k}}-\mu_{X,j}(P_{s_{k}})\right)\right)\geq d_{1-\beta/2}(\hat{\Sigma}^{s_{k}})\right)\geq 1-\beta/2$ (33)
yielding a contradiction of 31. Thus, we have:
$\liminf_{n\to\infty}\inf_{P\in\mathcal{P}}Pr_{P}\left(L^{n}\leq\Delta^{n}(P)\right)\geq 1-\beta/2$
with an analogous result holding for $U^{n}$. Applying a Bonferroni correction proves the lemma.
In addition, the following proposition demonstrates that $\rho_{1-(\alpha-\beta)}(\Delta^{n}(P),\Delta^{n}(P);\hat{\Sigma}^{n})$ is an asymptotically valid upper bound for the maximum, studentized deviation between all $\jmath$ in $\hat{J}_{R}$. Formally, we have the following lemma.
@@
Under assumptions B.1-B.2, the following uniform asymptotic validity condition applies to $\rho_{1-(\alpha-\beta)}$:
[…]
@@
By way of contradiction, let us assume that there exists a sequence of measures $P_{n_{s}}$ such that:
$\lim_{s\to\infty}Pr_{P_{n_{s}}}\left(\max_{j\in J}\left\{\frac{\sqrt{n_{s}}|\xi_{\tilde{S}_{Y}^{n_{s}},j}|}{\sqrt{\hat{\Sigma}_{Y,jj}^{n_{s}}}}\mathbb{1}\left(\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{\tilde{S}_{X}^{n_{s}},j}\geq\xi_{\tilde{S}_{X}^{n_{s}},j^{\prime}}+\Delta_{j^{\prime},j}\right)\in R\right)\right\}>\rho_{1-\alpha+\beta}(\Delta^{n_{s}},\Delta^{n_{s}};\hat{\Sigma}^{n_{s}})\right)=\eta>\alpha-\beta$
Equivalently:
$\lim_{s\to\infty}Pr_{P_{n_{s}}}\left(\max_{\jmath\in\hat{J}_{R}}\left\{\frac{\sqrt{n_{s}}\left|\xi_{\tilde{S}_{Y}^{n_{s}},\jmath}\right|}{\sqrt{\hat{\Sigma}_{Y,\jmath\jmath}^{n_{s}}}}\right\}>\rho_{1-\alpha+\beta}(\Delta^{n_{s}},\Delta^{n_{s}};\hat{\Sigma}^{n_{s}})\right)=\eta$
The same reasoning as in the proof of lemma B.2 provides a further subsequence of $n_{s}$ such that the above limit is less than or equal to $\alpha-\beta$, yielding a contradiction, and proving the lemma.
Finally, we can prove the following proposition:
Proposition B.1.
Given assumptions B.1-B.2, the two step confidence set is uniformly, asymptotically valid, such that:
$\liminf_{n\to\infty}\inf_{P\in\mathcal{P}}Pr_{P}\left(\left\{\mu_{Y,\jmath}^{n}(P)\right\}_{\jmath\in\hat{J}_{R}}\in CS_{1-\alpha;n}^{TS}\right)\geq 1-\alpha$ (34)
@@
Proof of Proposition B.1. First, we note that, by lemma B.2:
@@
$\limsup_{n\to\infty}\sup_{P\in\mathcal{P}}Pr_{P}\left(\left(\mu_{Y,\jmath}(P)\right)_{\jmath\in\hat{J}_{R}}\not\in CS_{1-\alpha;n}^{TS}\right)$
$\leq\limsup_{n\to\infty}\sup_{P\in\mathcal{P}}Pr_{P}\left(\left(\mu_{Y,\jmath}\left(P\right)\right)_{\jmath\in\hat{J}_{R}}\not\in CS_{1-\alpha;n}^{TS}\cap B^{n}\right)+Pr_{P}((B^{n})^{c})$
$\leq\limsup_{n\to\infty}\sup_{P\in\mathcal{P}}Pr_{P}\left(\left(\mu_{Y,\jmath}\left(P\right)\right)_{\jmath\in\hat{J}_{R}}\not\in CS_{1-\alpha;n}^{TS}\cap B^{n}\right)$
$\quad\quad\quad+\limsup_{n\to\infty}\sup_{P\in\mathcal{P}}Pr_{P}((B^{n})^{c})$
$\leq\limsup_{n\to\infty}\sup_{P\in\mathcal{P}}Pr_{P}\left(\left(\mu_{Y,\jmath}\left(P\right)\right)_{\jmath\in\hat{J}_{R}}\not\in CS_{1-\alpha;n}^{TS}\cap B^{n}\right)+\beta$
$\leq\limsup_{n\to\infty}\sup_{P\in\mathcal{P}}Pr_{P}\left(\bigcup_{\jmath\in\hat{J}_{R}}\left\{\frac{\sqrt{n}\left|\xi_{\tilde{S}_{Y}^{n},\jmath}\right|}{\sqrt{\hat{\Sigma}_{Y,\jmath\jmath}^{n}}}>\rho_{1-\alpha+\beta}\left(L^{n},U^{n};\hat{\Sigma}^{n}\right)\right\}\cap B^{n}\right)+\beta$
$\leq\limsup_{n\to\infty}\sup_{P\in\mathcal{P}}Pr_{P}\left(\bigcup_{\jmath\in\hat{J}_{R}}\left\{\frac{\sqrt{n}\left|\xi_{\tilde{S}_{Y}^{n},\jmath}\right|}{\sqrt{\hat{\Sigma}_{Y,\jmath\jmath}^{n}}}>\rho_{1-\alpha+\beta}\left(\Delta^{n},\Delta^{n};\hat{\Sigma}^{n}\right)\right\}\right)+\beta$
$\leq\limsup_{n\to\infty}\sup_{P\in\mathcal{P}}Pr_{P}\left(\max_{j\in J}\left\{\frac{\sqrt{n}|\xi_{\tilde{S}_{Y}^{n},j}|}{\sqrt{\hat{\Sigma}_{Y,jj}^{n}}}\mathbb{1}\left(\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{\tilde{S}_{X}^{n},j}\geq\xi_{\tilde{S}_{X}^{n},j^{\prime}}+\Delta_{j^{\prime},j}\right)\in R\right)\right\}>\rho_{1-\alpha+\beta}(\Delta^{n},\Delta^{n};\hat{\Sigma}^{n})\right)+\beta$
@@
where the final inequality holds by lemma B.3.

### v4

@@ Appendix B Uniform Asymptotic Validity and Proofs
Let us recall the following inequality:
$(Z_{1}+Z_{2})^{2}\mathbb{1}\left(|Z_{1}+Z_{2}|>\lambda\right)\leq 2\left(Z_{1}^{2}\mathbb{1}\left(|Z_{1}|>\lambda/2\right)+Z_{2}^{2}\mathbb{1}\left(|Z_{2}|>\lambda/2\right)\right)$
For any $j$ and $j^{\prime}$, such that $j\neq j^{\prime}$ and such that $\text{var}_{jj^{\prime}}(P)\neq 0$, we take $Z_{1}=\tilde{W}_{1,j}-\mu_{W,j}(P)$ and $Z_{2}=\tilde{W}_{1,j^{\prime}}-\mu_{W,j^{\prime}}(P)$. We choose $\lambda/2$ to satisfy assumption B.1 with $\text{var}_{jj^{\prime}}(P)\varepsilon/2$ as the upper bound in (33). Taking $K:=\lambda\sqrt{\text{var}_{jj^{\prime}}(P)}$ yields the lemma.
We also provide the following technical lemma. Let $R_{n}(\tilde{S}_{W}^{n},P_{n})$ be some function of $\tilde{S}_{W}^{n}$, with the cumulative distribution function $J_{n}(x,P_{n})$ under $P_{n}$. Let $\tilde{P}_{n}$ be some sequence of random distributions. The following lemma holds:
@@
Suppose, for any $\varepsilon>0$, that the sequence of $\tilde{P}_{n}$ satisfies the following:
$\liminf_{n\to\infty}\inf_{P\in\mathcal{P}}Pr_{P}\left(\sup_{x\in\mathbb{R}}\left|J_{n}(x,\tilde{P}_{n})-J_{n}(x,P)\right|\leq\varepsilon\right)=1$ (34)
Then, for any $0\leq\alpha_{1}$, $0\leq\alpha_{2}$ such that $0\leq\alpha_{1}+\alpha_{2}<1$, the following holds:
$\liminf_{n\to\infty}\inf_{P\in\mathcal{P}}Pr_{P}\left(J_{n}^{-1}(\alpha_{1},\tilde{P}_{n})\leq R_{n}(\tilde{S}_{W}^{n},P_{n})\leq J_{n}^{-1}(1-\alpha_{2},\tilde{P}_{n})\right)\geq 1-\alpha_{1}-\alpha_{2}$ (35)
@@
Let us fix arbitrary $\eta>0$. For any $n$ sufficiently large, we have:
$1-\frac{\eta}{2}\leq\inf_{P\in\mathcal{P}}Pr_{P}\left(\sup_{x\in\mathbb{R}}\left|J_{n}(x,\tilde{P}_{n})-J_{n}(x,P)\right|\leq\frac{\eta}{2}\right)$
This implies, by part viii of lemma A.1 of Romano and Shaikh (2012), that:
$\inf_{P\in\mathcal{P}}Pr_{P}\left(J_{n}^{-1}(\alpha_{1},\tilde{P}_{n})\leq R_{n}(\tilde{S}_{W}^{n},P_{n})\leq J_{n}^{-1}(1-\alpha_{2},\tilde{P}_{n})\right)\geq 1-\alpha_{1}-\alpha_{2}-\eta$
Finally, because $\eta$ was arbitrary, we find that (35) holds, thus proving the lemma.
The following lemma concerns the asymptotic properties of $L^{n}$, and namely, whether $L^{n}$ satisfies a uniform, asymptotic validity condition as a lower bound of $\Delta^{n}_{j^{\prime}j}\equiv\Delta^{n}_{j^{\prime}j}(P):=\mu_{X,j^{\prime}}(P)-\mu_{X,j}(P)$. We naturally define $L^{n}$ and $U^{n}$ as follows:
$L_{j^{\prime}j}^{n}$ $=\tilde{S}_{X,j^{\prime}}^{n}-\tilde{S}_{X,j}^{n}-d_{1-\beta/2}(\hat{\Sigma}^{n})\sqrt{\hat{\text{var}}^{n}_{j^{\prime}j}}$
$U_{j^{\prime}j}^{n}$ $=\tilde{S}_{X,j^{\prime}}^{n}-\tilde{S}_{X,j}^{n}+d_{1-\beta/2}(\hat{\Sigma}^{n})\sqrt{\hat{\text{var}}^{n}_{j^{\prime}j}}$
where $\hat{\text{var}}^{n}_{j^{\prime}j}=\hat{\Sigma}^{n}_{j^{\prime}j^{\prime}}+\hat{\Sigma}^{n}_{jj}-2\hat{\Sigma}^{n}_{j^{\prime}j}$, with an analogous definition for $\text{var}_{j^{\prime}j}(P)$.
Lemma B.4.
Let us assume that assumption B.1 holds. Under these conditions, it follows that $L^{n}$ is a uniformly, asymptotically valid lower bound for $\Delta^{n}(P)$. In particular:
@@
We will prove that $L^{n}$ is a uniformly valid lower bound for $\Delta^{n}(P)$ at level $1-\beta/2$. The joint result in (36) follows by a union bound. Before proceeding, we notice that, by lemma B.2 and lemma s.6.1 of Romano and Shaikh (2012), we obtain following uniform consistency condition on the $\hat{\text{var}}_{j^{\prime}j}^{n}$:
$\limsup_{n\to\infty}\sup_{P\in\mathcal{P}}Pr_{P}\left(\left|\frac{n\hat{\text{var}}_{jj^{\prime}}^{n}}{\text{var}_{jj^{\prime}}(P)}-1\right|>\varepsilon\right)=0\quad\text{for $\varepsilon>0$, $j,j^{\prime}\in J$ such that $\text{var}_{jj^{\prime}}(P)>0$}$ (37)
The remainder of our proof will use the results of lemma B.3. In particular, we define $R_{n}(\tilde{S}_{W}^{n},P)$ as follows:
$\max_{j,j^{\prime}\in J,\text{var}_{j^{\prime}j}(P)\neq 0}\frac{\sqrt{n}\left((\tilde{S}_{X,j}^{n}-\mu_{X,j}(P))-(\tilde{S}_{X,j^{\prime}}^{n}-\mu_{X,j^{\prime}}(P))\right)}{\sqrt{\text{var}_{j^{\prime}j}(P)}}$ (38)
Noting that the event that $R_{n}(\tilde{S}_{W}^{n},P)\leq x$ holds if and only if $\sqrt{n}(\tilde{S}_{W}^{n}-\mu_{W}(P))$ lies in a convex set $S$ contained in $\mathcal{C}$, where we define $\mathcal{C}$ in lemma B.1, the same lemma gives us that:
$\sup_{P\in\mathcal{P}}\sup_{x\in\mathbb{R}}\left|J_{n}(x,P)-J_{n}(x,\Phi_{{\Sigma(P)}})\right|=o(1)$
Moreover, the continuous mapping theorem, the uniform consistency result in (37), lemma s.7.1 in Romano and Shaikh (2008), and Polya’s theorem provide that, for any $\varepsilon>0$:
$\liminf_{n\to\infty}\inf_{P\in\mathcal{P}}Pr_{P}\left(\sup_{x\in\mathbb{R}}\left|J_{n}(x,\Phi_{{\hat{\Sigma}^{n}}})-J_{n}(x,\Phi_{{\Sigma(P)}})\right|\leq\varepsilon\right)=1$
or equivalently:
$\liminf_{n\to\infty}\inf_{P\in\mathcal{P}}Pr_{P}\left(\sup_{x\in\mathbb{R}}\left|J_{n}(x,\tilde{P}_{n})-J_{n}(x,\Phi_{{\Sigma(P)}})\right|\leq\varepsilon\right)=1$
where $\tilde{P}_{n}$ corresponds to the distribution of a gaussian with mean $\mu_{W}(P)$ and covariance $\hat{\Sigma}^{n}$. Thus, we can verify:
$\liminf_{n\to\infty}\inf_{P\in\mathcal{P}}Pr_{P}\left(\sup_{x\in\mathbb{R}}\left|J_{n}(x,\tilde{P}_{n})-J_{n}(x,P)\right|\leq\varepsilon\right)=1$
Applying lemma B.3 gives us that:
$\liminf_{n\to\infty}\inf_{P\in\mathcal{P}}Pr_{P}\left(J_{n}^{-1}(0,\tilde{P}_{n})\leq R_{n}(\tilde{S}_{W}^{n},P)\leq J_{n}^{-1}\left(1-\frac{\beta}{2},\tilde{P}_{n}\right)\right)\geq 1-\frac{\beta}{2}$
We can note that $J_{n}^{-1}(0,\tilde{P}_{n})=0$, while $J_{n}^{-1}(1-\beta/2,\tilde{P}_{n})=d_{1-\beta/2}(\hat{\Sigma}^{n})$. Consequently, algebraic manipulation of $R_{n}(\tilde{S}_{W}^{n},P)$ as defined in (38) and an application of the uniform consistency result in (37) jointly imply that (36) holds, proving the lemma.
In addition, the following proposition demonstrates that $\rho_{1-(\alpha-\beta)}(\Delta^{n}(P),\Delta^{n}(P);\hat{\Sigma}^{n})$ is an asymptotically valid upper bound for the maximum, studentized deviation between all $\hat{\jmath}$ in $\hat{J}_{R}$. Formally, we have the following lemma.
Lemma B.5.
Under assumption B.1, the following uniform asymptotic validity condition applies to $\rho_{1-(\alpha-\beta)}$:
$\limsup_{n\to\infty}\sup_{P\in\mathcal{P}}Pr_{P}\left(\max_{j\in J}\left\{\frac{|\xi_{\tilde{S}_{Y}^{n},j}|}{\sqrt{\hat{\Sigma}_{Y,jj}^{n}}}\mathbb{1}\left(\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{\tilde{S}_{X}^{n},j}\geq\xi_{\tilde{S}_{X}^{n},j^{\prime}}+\Delta_{j^{\prime}j}\right)\in R\right)\right\}>\rho_{1-\alpha+\beta}(\Delta(P),\Delta(P);\hat{\Sigma}^{n})\right)\leq\alpha-\beta$
Proof.
Our proof proceeds much as the proof of lemma B.4. Indeed, we define $R_{n}(\tilde{S}_{W}^{n},P)$ as follows:
$\max_{j\in J}\frac{\left|\sqrt{n}\left(\tilde{S}_{Y,j}^{n}-\mu_{Y,j}(P)\right)\right|}{\sqrt{\Sigma_{Y,jj}(P)}}\mathbb{1}\left(\sum_{j^{\prime}\in J}\mathbb{1}\left(\sqrt{n}\left(\tilde{S}_{X,j}^{n}-\mu_{X,j}(P)\right)-\sqrt{n}\left(\tilde{S}_{X,j^{\prime}}^{n}-\mu_{X,j^{\prime}}(P)\right)\geq\sqrt{n}\Delta_{j^{\prime}j}(P)\right)\in R\right)$
The indicator in the above expression is equivalent to the indicator $\mathbb{1}(j\in\hat{J}_{R;n})$, where $\hat{J}_{R;n}=\hat{J}_{R}(\tilde{S}_{X}^{n})$. Thus, we seek to show that the event:
$\max_{j\in\hat{J}_{R;n}}\frac{\left|\sqrt{n}\left(\tilde{S}_{Y,j}^{n}-\mu_{Y,j}(P)\right)\right|}{\sqrt{\Sigma_{Y,jj}(P)}}\leq x$ (40)
lies in $\mathcal{C}$ as defined in assumption B.1. The above event can be rewritten as:
$A(x):=\bigcup_{J_{c}\in 2^{J}}\left(\left\{\max_{j\in J_{c}}\frac{\left|\sqrt{n}\left(\tilde{S}_{Y,j}^{n}-\mu_{Y,j}(P)\right)\right|}{\sqrt{\Sigma_{Y,jj}(P)}}\leq x\right\}\bigcap\left\{\hat{J}_{R;n}=J_{c}\right\}\right)$
We may notice that:
$\partial A(x)\subseteq\bigcup_{J_{c}\in 2^{J}}\left\{\max_{j\in J_{c}}\frac{\left|\sqrt{n}\left(\tilde{S}_{Y,j}^{n}-\mu_{Y,j}(P)\right)\right|}{\sqrt{\Sigma_{Y,jj}(P)}}=x\right\}=:S$
The right hand side of $S$ is a subset of the boundary of a particular hyperrectangle, and thus is such that $\Phi_{V}(S)=0$ for any $V$ as described in lemma B.1. Now, the following hold:
$\sup_{P\in\mathcal{P}}\sup_{x\in\mathbb{R}}\left|J_{n}(x,P)-J_{n}(x,\Phi_{{\Sigma(P)}})\right|$ $=o(1)$
$\liminf_{n\to\infty}\inf_{P\in\mathcal{P}}Pr_{P}\left(\sup_{x\in\mathbb{R}}\left|J_{n}(x,\tilde{P}_{n})-J_{n}(x,\Phi_{{\Sigma(P)}})\right|\leq\varepsilon\right)$ $=1$
$\liminf_{n\to\infty}\inf_{P\in\mathcal{P}}Pr_{P}\left(\sup_{x\in\mathbb{R}}\left|J_{n}(x,\tilde{P}_{n})-J_{n}(x,P)\right|\leq\varepsilon\right)$ $=1$
$\liminf_{n\to\infty}\inf_{P\in\mathcal{P}}Pr_{P}\left(J_{n}^{-1}(0,\tilde{P}_{n})\leq R_{n}(\tilde{S}_{W}^{n},P)\leq J_{n}^{-1}\left(1-\alpha+\beta,\tilde{P}_{n}\right)\right)$ $\geq 1-\alpha+\beta$ (41)
where $\tilde{P}_{n}$ denotes the distribution of a gaussian with mean $\mu_{W}(P)$ and variance $\hat{\Sigma}^{n}$. The first equality holds by an application of lemma B.1. The second equality holds by lemmas s.6.1 and s.7.1 in Romano and Shaikh (2012), and Polya’s theorem. The third equality is a consequence of the first two equalities, and the final equality is a consequence of B.3. Finally, we may notice that $J_{n}^{-1}(0,\tilde{P}_{n})=0$, and $J_{n}^{-1}\left(1-\alpha+\beta\right)=\rho_{1-\alpha+\beta}(\Delta(P),\Delta(P);\hat{\Sigma}^{n})$ by construction. We also notice that assumption B.1 implies that for any $\varepsilon>0$:
$\limsup_{n\to\infty}\sup_{P\in\mathcal{P}}Pr_{P}\left(\left|\frac{n\hat{\Sigma}_{Y,jj}^{n}}{\Sigma_{Y,jj}(P)}-1\right|>\varepsilon\right)=0$
by lemma s.6.1 of Romano and Shaikh (2012). This result, along with (41) gives that (B.5) holds, thus proving the lemma.
Finally, we can prove proposition 3.2:
Proof.
Proof of Proposition 3.2. First, we note that, by lemma B.4:
@@
$\limsup_{n\to\infty}\sup_{P\in\mathcal{P}}Pr_{P}\left(\left(\mu_{Y,\hat{\jmath}}(P)\right)_{\hat{\jmath}\in\hat{J}_{R}}\not\in CS_{1-\alpha;n}^{TS}\right)$
$\leq\limsup_{n\to\infty}\sup_{P\in\mathcal{P}}Pr_{P}\left(\left(\mu_{Y,\hat{\jmath}}\left(P\right)\right)_{\hat{\jmath}\in\hat{J}_{R}}\not\in CS_{1-\alpha;n}^{TS}\cap B^{n}\right)+Pr_{P}((B^{n})^{c})$
$\leq\limsup_{n\to\infty}\sup_{P\in\mathcal{P}}Pr_{P}\left(\left(\mu_{Y,\hat{\jmath}}\left(P\right)\right)_{\hat{\jmath}\in\hat{J}_{R}}\not\in CS_{1-\alpha;n}^{TS}\cap B^{n}\right)+\limsup_{n\to\infty}\sup_{P\in\mathcal{P}}Pr_{P}((B^{n})^{c})$
$\leq\limsup_{n\to\infty}\sup_{P\in\mathcal{P}}Pr_{P}\left(\left(\mu_{Y,\hat{\jmath}}\left(P\right)\right)_{\hat{\jmath}\in\hat{J}_{R}}\not\in CS_{1-\alpha;n}^{TS}\cap B^{n}\right)+\beta$
$\leq\limsup_{n\to\infty}\sup_{P\in\mathcal{P}}Pr_{P}\left(\bigcup_{\hat{\jmath}\in\hat{J}_{R}}\left\{\frac{\sqrt{n}\left|\xi_{\tilde{S}_{Y}^{n},\hat{\jmath}}\right|}{\sqrt{\hat{\Sigma}_{Y,\hat{\jmath}\hat{\jmath}}^{n}}}>\rho_{1-\alpha+\beta}\left(L^{n},U^{n};\hat{\Sigma}^{n}\right)\right\}\cap B^{n}\right)+\beta$
$\leq\limsup_{n\to\infty}\sup_{P\in\mathcal{P}}Pr_{P}\left(\bigcup_{\hat{\jmath}\in\hat{J}_{R}}\left\{\frac{\sqrt{n}\left|\xi_{\tilde{S}_{Y}^{n},\hat{\jmath}}\right|}{\sqrt{\hat{\Sigma}_{Y,\hat{\jmath}\hat{\jmath}}^{n}}}>\rho_{1-\alpha+\beta}\left(\Delta^{n},\Delta^{n};\hat{\Sigma}^{n}\right)\right\}\right)+\beta$
$\leq\limsup_{n\to\infty}\sup_{P\in\mathcal{P}}Pr_{P}\left(\max_{j\in J}\left\{\frac{\sqrt{n}|\xi_{\tilde{S}_{Y}^{n},j}|}{\sqrt{\hat{\Sigma}_{Y,jj}^{n}}}\mathbb{1}\left(\sum_{j^{\prime}\in J}\mathbb{1}\left(\xi_{\tilde{S}_{X}^{n},j}\geq\xi_{\tilde{S}_{X}^{n},j^{\prime}}+\Delta_{j^{\prime}j}\right)\in R\right)\right\}>\rho_{1-\alpha+\beta}(\Delta^{n},\Delta^{n};\hat{\Sigma}^{n})\right)+\beta$
@@
where the third inequality follows by lemma B.4 and the final inequality holds by lemma B.5.